# A Calculus of Overlays

### v2

@@ C.4. Existence of the Least Fixed Point
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
@@ D.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the overlay tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting overlay tree has a specific shape:
@@
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the overlay tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different overlay trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of overlay-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
@@ D.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
@@
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overlays}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one overlay at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{overlay}})$ in $\mathtt{supers}$.
@@
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overlays}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
@@ D.4. Substitution Lemma
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
@@ D.5. Convergence Preservation
That is, following one $\mathrm{result}$ projection in the tail call’s overlay tree yields the same properties as the root of the reduct’s overlay tree.
@@ D.7. Full Abstraction
Adequacy relates a single term’s head normal form to its overlay-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition D.5 (Overlay-contextual equivalence).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
@@
This definition observes only convergence (a single bit), but quantifying over all contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior.
@@
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the overlay tree $\mathcal{T}(C[M])$ is determined by the overlay tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993). The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.

### v4

@@ C.4. Existence of the Least Fixed Point
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
@@ D.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the mixin tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
@@
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of inheritance-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
@@ D.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
@@
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
@@
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
@@ D.4. Substitution Lemma
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
@@ D.5. Convergence Preservation
That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
@@ D.7. Full Abstraction
Adequacy relates a single term’s head normal form to its inheritance-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition D.5 ($\lambda$-contextual equivalence under $\mathcal{T}$).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
@@
The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts. This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
@@
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993). The mixin tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside mixin trees.